OpenStudy (itsstribe):

The perimeter of a rectangle is 90 centimeters. The width of the rectangle can be represented by (x+2), and its length can be represented by (3x-5). What are the dimensions of this rectangle in centimeters?

OpenStudy (erak):

Try setting up a perimeter equation for a rectangle like so: P = 2w + 2l and sub in those binomials

OpenStudy (itsstribe):

So can you keep showing me through

OpenStudy (itsstribe):

I can't find how to put it togethor

OpenStudy (erak):

okay so w = (x+2) and l = (3x-5) right, once you put it into the equation it will look like P = 2(x+2) + 2(3x-5)

OpenStudy (itsstribe):

i tried using calculator and got 8x-6 for p

OpenStudy (erak):

Ah they also give you P so it's 90 = 2(x+2) + 2(3x-5)

OpenStudy (itsstribe):

So it would be b?

OpenStudy (erak):

Try solving 90 = 8x-6

OpenStudy (itsstribe):

Is b not the answer? https://gyazo.com/564e76ea8194f595851d84037e0b3d56

OpenStudy (erak):

Not quite, try solving for x in the equation I put and then plug that back into w = (x+2) and l = (3x-5)

OpenStudy (itsstribe):

x=12

OpenStudy (itsstribe):

is the answer C?

OpenStudy (erak):

yeah